OpenStudy (anonymous):

If the monopolist's demand curve is P = 50 − 10Q, then marginal revenue is zero when Q is equal to

OpenStudy (anonymous):

P=50-Q MR is twice as steep as the demand curve. Therefore P=50-2Q MR=0 when Q=2.5